The formula for the approximate yield of an investment can look intimidating, but its really just a function of three things: (1) average current income, (2) average capital gains, and (3) the average value of the investment. Based on the information in the table, compute each of these values for the two stocks over a 3-year periad and enter the values into the bottom half of the table. Stock 1 Expected average anual dividends (2012-2014) 140 Current stock price Expected future stock price (2014) Stock 2 $3.10 $115 $151 $60 578 A-Z Average current income (CI) Average capital gains (CG) Average value of the investment (VI) Next, derive the correct fomula for approximate yield by correctly arranging these three variables in the cquation that follows. CI+Avg VI Approximate Yield Using this formula, you can see that the approximate yield for Stock 1 is 0.09%-and the approximate yield for Stock 2 is 11.34% True or False: If both investments carry the same rate of risk, Stock 2 is a better investment than Stock 1 True False 3:36 PM 210/2018
Answer #1

Approximate yield is calculated as follows -

Approximate yield = (Avg CI + Avg CG) / Avg VI

Particulars Stock 1 Stock 2
Avg CI (Dividend income) $1.40 × 3/ 3 = $1.40 $3.10 × 3/ 3 = $3.10
Avg CG ($78 - $60)/ 3 = $6 ($151 - $115)/ 3 = $12
Avg VI ($60+$78)/ 2 = $69 ($151 + $115)/ 2 = $133

Stock 1 yield = ($1.40 + $6)/ $69 = 0.107246 or 10.72%

Stock 2 yield = ($3.10 + $12)/ $133 = 0.113534 or 11.35%

Yes, if both have same risk, then stock 2 is better as it has better yield.